OpenStudy (anonymous):

a line with a slope of -2 passes thru the coordinate (2,-1) and also passes thru the (2.y) a) determine the value for y b) determine another point on the line

OpenStudy (anonymous):

y=mx+c where m is the slope. simples

OpenStudy (anonymous):

change in y --------- change in x y--1 ---- 2-2 y+1=0(-2) y+1=0 y=-1